# April 2025

### v1

@@ ABSTRACT
Is there a Hierarchy Problem? If so, what, exactly, is the problem? Almost every theorist has a personal answer to these questions. In this article, I give my answers. I will explain that the Hierarchy Problem is not a formal problem but rather our ignorance of a crucial physics explanation – the explanation of the nature of the Higgs boson. Without the solution to this problem, we cannot make progress on the major questions of our field.
@@ 1 Introduction
Today, the “Hierarchy Problem” is the question in particle physics about which there is the least consensus. There are many conceptions of this problem and the related question of the “naturalness” of the parameters of the Standard Model. I doubt that — absent any illuminating experimental discovery — a new theoretical paper could meet the stated goal of this special issue of “clarifying common misconceptions”. Still, I am glad to have this opportunity to state my personal opinions on what the Hierarchy Problem really asks and offer my own route to a solution. As to whether I am dissolving confusion and simply adding to it, I leave to the reader to judge.
@@ 2 Is there a Hierarchy Problem?
But one should also consider the physical meaning of this subtraction. The massless particles implied by (3) do not exist. If any higher-mass particle of mass $M$ contributes to the loop, it will, in dimensional regularization, give a contribution of the structure of one of the terms in (2) with the $\Lambda^{2}$ replaced by $M^{2}$. This includes particles assocated with the completion of the SM to include quantum gravity. Fluctuations of the $t$, $\Phi$, and $W$, $Z$ quantum fields at high momenta certainly do exist. Even dimensional regularization tells us that these fluctuations contribute to the renormalization group $\beta$ function of the $\lambda$ parameter. In fact, they contribute to the prediction of the SM that its vacuum is unstable [2, 3]. So their effect is not so easily dismissed [4].
@@
Before going further, though, I would like to call attention to a benefit of taking the fluctuations of the top quark field at large momentum seriously. Consider the vacuum of the top quark quantum field as represented by a Dirac sea. It is obvious that generating a HIggs field vacuum expectation value that gives the top quark a mass opens a mass gap in the spectrum that lowers the energy of the filled Dirac sea, which is the same as lowering the energy of the vacuum. I have found this argument helpful in clarifying the mystery of EWSB for physics department colloquium audiences. Why not take it seriously? The idea that the largeness of the top quark Yukawa coupling relative to other dimensionless couplings of the SM is the driver of EWSB is very enticing. This was first noticed in the context of supersymmetric grand unified theories [5, 6, 7], but it applies in a very wide variety of explicit models of EWSB, as I will discuss below.
@@ 3 The Wilsonian point of view
As a student of Wilson, I am thoroughly infected with that philosophy, for better or worse. Spontaneous breaking of symmetry is found in a great variety of condensed matter systems — magnets, superconductors, liquid crystals, and more. The condensed matter context forces us to think about symmetry-breaking in these systems in a different way. Condensed matter systems are made of atoms, and the laws of atomic physics have been fixed since discovery of quantum mechanics. This means that the explanations for these examples of spontaneous symmetry breaking must be derivable from those laws. This makes the study of phase transitions in condensed matter a fascinating subject, with many surprising mechanisms leading to the observed nontrivial ground states [10]. Thinking in this way get us away from simplifying the hierarchy program to the explanation of a large ratio of mass scales.
@@ 4 The Hierarchy Problem before the LHC Results
For me, though, supersymmetry had a different attraction. Supersymmetry provides a quantitative physical mechanism by which the top quark Yukawa coupling can provide a physical mechanism for EWSB. In supersymmetric theories, the logarithmic renormalization of soft supersymmetry breaking scalar mass terms by Yukawa couplings is negative. So, if supersymmetry is spontaneously broken, as it must be to provide a realistic theory, then the renormalization of the symmetry-breaking mass terms will lead to a vacuum instability for one or more scalar fields. There are many, many new scalar fields in a supersymmetric extension of the SM, leading to many possibilities for symmetry breaking. But, due to the large size of the top quark Yukawa coupling and some convenient group theory factors, it is the Higgs doublet scalars for which this effect is largest, leading to the exactly the pattern of EWSB found in the SM [5, 6, 7]. This mechanism is highlighted in my 2006 TASI lectures on supersymmetry [13].
@@ 5 The “Post-Naturalness Era”
Excuse me that I am very cool to these ideas. Particle theorists often view scalars as part of the general equipment of nature, to be added at will to any theory. As I have already stated above, I believe that scalar fields in nature should have a purpose. If they obtain vacuum expectation values, those expectation values should be tied to the masses of particles or to another dimensionful reference point. It is possible that that the vacuum expectation value of the Higgs field could be chosen randomly. But it would be much more pleasing to explain this value in an underly physical picture.
@@ 6 Three Hierarchy Problems
The second is what one might call the “Problem of Scalars”. This is the flip side of the idea that one should add scalars to a model to give it more interesting dynamics. Scalars have dimension 1 and can couple to other fields almost without restriction. Even extending the Higgs sector of the SM to two Higgs doublets extends the parameter set to 4 mass parameters and 10 quartic parameters, counting possibly complex coefficients as two real parameters.
@@
It has been all downhill since then. Despite the efforts of many of the leading theorists of the my generation — including Jogesh Pati, Savas Dimopoulos, Lawrence Hall, Graham Ross, and Yossi Nir — no one would say today, “Give me one more decimal place on the CKM parameters and this problem will be solved.” The explanation for the neutrino masses, which involve a $3\times 3$ complex seesaw mass matrix in additional to a Yukawa matrix, is even further away.
In addressing this problem, it would be tremendously illuminating to discover the organizing principle for the scalar couplings. Do they come from compositeness the Higgs boson? from the mixing of the lightest Higgs field with heavier partners? from a hierarchy of couplings generated by perturbation theory? An aspect of quantum field theory that we know little about today is the dynamics and possible composite particles of chiral gauge theories. Maybe this is a place that we will find clues to these questions.
@@ 8 The Hierarchy Problem and future high-energy colliders
Today, 10 TeV pCM colliders are still out of reach. In the next decades, the particle physics community must put serious effort into the development of new tools and technologies to reach those energies. In parallel, theorists must work to define the experimental program that these accelerators will carry out. We cannot do this without confronting the Hierarch Problem and bringing new ideas to its solution. This is a responsibility that the particle theory community must address.
@@ ACKNOWLEDGEMENTS
This work was supported by the US Department of Energy under contract DE–AC02–76SF00515. I am grateful to Nathaniel Craig and David Gross for a visit to the the Kavli Institute for Theoretical Physics that allowed me to preview these ideas before a skeptical and discerning audience. That visit was supported by National Science Foundation grant NSF PHY-2309135 to the KITP.

### v3

# SLAC-PUB-250501 May 2025

@@ ABSTRACT
Is there a Hierarchy Problem? If so, what, exactly, is the problem? Almost every theorist has a personal answer to these questions. In this article, I give my answer. I will explain that the Hierarchy Problem is not a formal problem but rather our ignorance of a crucial physics explanation – the explanation of the nature of the Higgs boson. Without the solution to this problem, we cannot make progress on the major questions of our field.
@@ 1 Introduction
Today, the “Hierarchy Problem” is the question in particle physics about which there is the least consensus. There are many conceptions of this problem and the related question of the “naturalness” of the parameters of the Standard Model. I doubt that — absent any illuminating experimental discovery — a new theoretical paper could meet the stated goal of this special issue of “clarifying common misconceptions”. Still, I am glad to have this opportunity to state my personal opinions on what the Hierarchy Problem really asks and to offer my own route to a solution. As to whether I am dissolving confusion and simply adding to it, I leave to the reader to judge.
@@ 2 Is there a Hierarchy Problem?
But one should also consider the physical meaning of this subtraction. The massless particles implied by (3) do not exist. If any higher-mass particle of mass $M$ contributes to the loop, it will, in dimensional regularization, give a contribution of the structure of one of the terms in (2) with the $\Lambda^{2}$ replaced by $M^{2}$. This includes particles associated with the completion of the SM to include quantum gravity. The specific example of a contribution from the heavy neutrino mass $M_{R}^{2}$ in the seesaw mechanism of neutrino mass generation is worked out in [2]. Fluctuations of the $t$, $\Phi$, and $W$, $Z$ quantum fields at high momenta certainly do exist. Even dimensional regularization tells us that these fluctuations contribute to the renormalization group $\beta$ function of the $\lambda$ parameter. In fact, they contribute to the prediction of the SM that its vacuum is unstable [3, 4]. So their effect is not so easily dismissed [5].
@@
Before going further, though, I would like to call attention to a benefit of taking the fluctuations of the top quark field at large momentum seriously. Consider the vacuum of the top quark quantum field as represented by a Dirac sea. It is obvious that generating a Higgs field vacuum expectation value that gives the top quark a mass opens a mass gap in the spectrum that lowers the energy of the filled Dirac sea, which is the same as lowering the energy of the vacuum. I have found this argument helpful in clarifying the mystery of EWSB for physics department colloquium audiences. Why not take it seriously? The idea that the largeness of the top quark Yukawa coupling relative to other dimensionless couplings of the SM is the driver of EWSB is very enticing. This was first noticed in the context of supersymmetric grand unified theories [6, 7, 8, 9], but it applies in a very wide variety of explicit models of EWSB, as I will discuss below.
@@ 3 The Wilsonian point of view
As a student of Wilson, I am thoroughly infected with that philosophy, for better or worse. Spontaneous breaking of symmetry is found in a great variety of condensed matter systems — magnets, superconductors, liquid crystals, and more. The condensed matter context forces us to think about symmetry-breaking in these systems in a different way. Condensed matter systems are made of atoms, and the laws of atomic physics have been fixed since discovery of quantum mechanics. This means that the explanations for these examples of spontaneous symmetry breaking must be derivable from those laws. This makes the study of phase transitions in condensed matter a fascinating subject, with many surprising mechanisms leading to the observed nontrivial ground states [12]. Thinking in this way gets us away from simplifying the hierarchy program to the explanation of a large ratio of mass scales.
@@ 4 The Hierarchy Problem before the LHC Results
For me, though, supersymmetry had a different attraction. Supersymmetry provides a quantitative physical mechanism by which the top quark Yukawa coupling can provide a physical mechanism for EWSB. In supersymmetric theories, the logarithmic renormalization of soft supersymmetry breaking scalar mass terms by Yukawa couplings is negative. So, if supersymmetry is spontaneously broken, as it must be to provide a realistic theory, then the renormalization of the symmetry-breaking mass terms will lead to a vacuum instability for one or more scalar fields. There are many, many new scalar fields in a supersymmetric extension of the SM, leading to many possibilities for symmetry breaking. But, due to the large size of the top quark Yukawa coupling and some convenient group theory factors, it is the Higgs doublet scalars for which this effect is largest, leading to exactly the pattern of EWSB found in the SM [6, 7, 8, 9]. This mechanism is highlighted in my 2006 TASI lectures on supersymmetry [15].
@@ 5 The “Post-Naturalness Era”
Excuse me that I am very cool to these ideas. Particle theorists often view scalars as part of the general equipment of nature, to be added at will to any theory. As I have already stated above, I believe that scalar fields in nature should have a purpose. If they obtain vacuum expectation values, those expectation values should be tied to the masses of particles or to another dimensionful reference point. It is possible that the vacuum expectation value of the Higgs field could be chosen randomly. But it would be much more pleasing to explain this value in an underly physical picture.
@@ 6 Three Hierarchy Problems
The second is what one might call the “Problem of Scalars”. This is the flip side of the idea that one should add scalars to a model to give it more interesting dynamics. Scalars have dimension 1 and can couple to other fields almost without restriction. Even extending the Higgs sector of the SM to two Higgs doublets extends the parameter set to 4 mass parameters and 10 quartic parameters, counting possibly complex coefficients as two real parameters (reducible to a total of 11 after field redefinitions).
@@
It has been all downhill since then. Despite the efforts of many of the leading theorists of my generation — including Jogesh Pati, Savas Dimopoulos, Lawrence Hall, Graham Ross, and Yossi Nir — no one would say today, “Give me one more decimal place on the CKM parameters and this problem will be solved.” The explanation for the neutrino masses, which involve a $3\times 3$ complex seesaw mass matrix in additional to a Yukawa matrix, is even further away.
In addressing this problem, it would be tremendously illuminating to discover the organizing principle for the scalar couplings. Do they come from compositeness of the Higgs boson? from the mixing of the lightest Higgs field with heavier partners? from a hierarchy of couplings generated by perturbation theory? An aspect of quantum field theory that we know little about today is the dynamics and possible composite particles of chiral gauge theories. Maybe this is a place that we will find clues to these questions.
@@ 8 The Hierarchy Problem and future high-energy colliders
Today, 10 TeV pCM colliders are still out of reach. In the next decades, the particle physics community must put serious effort into the development of new tools and technologies to reach those energies. In parallel, theorists must work to define the experimental program that these accelerators will carry out. We cannot do this without confronting the Hierarchy Problem and bringing new ideas to its solution. This is a responsibility that the particle theory community must address.
@@ ACKNOWLEDGEMENTS
This work was supported by the US Department of Energy under contract DE–AC02–76SF00515. I am grateful to Nathaniel Craig and David Gross for a visit to the Kavli Institute for Theoretical Physics that allowed me to preview these ideas before a skeptical and discerning audience. That visit was supported by National Science Foundation grant NSF PHY-2309135 to the KITP.